The Kb of ammonia is 1.76 × 10-5. The pH of a buffer prepared by combining 45.0 mL of 0.180 M ammonia and 40.0 mL of 0.205 M ammonium nitrate is __________.


1
Expert's answer
2021-11-22T09:17:04-0500

Solution:

ammonia = NH3

ammonium nitrate = NH4NO3


Recalculate the molar concentration of each species:

VSolution = VNH3 + VNH4NO3 = 45.0 mL + 40.0 mL = 85.0 mL

CM(NH3) = Co(NH3) × (VNH3 / VSolution) = (0.180 M) × (45.0 mL / 85.0 mL) = 0.0953 M

[NH3] = 0.0953 M

CM(NH4NO3) = Co(NH4NO3) × (VNH4NO3 / VSolution) = (0.205 M) × (40.0 mL / 85.0 mL) = 0.0965 M

[NH4+] = 0.0965 M


To calculate the pH of a given buffer, you need to use the Henderson-Hasselbalch equation for basic buffers: pH = 14 - (pKb + log([B+]/[BOH])


[B+] = [NH4+] = 0.0965 M

[BOH] = [NH3] = 0.0953 M

pKb = -log(Kb) = -log(1.76×10-5) = 4.75

Hence,

pH = 14 - (4.75 + log(0.0965 / 0.0953) = 9.24

pH = 9.24


Answer: 9.24